### v2

# LLM Pruning and Distillation in Practice: The Minitron Approach

Abstract
Abstract: We present a comprehensive report on compressing the Llama 3.1 8B and Mistral NeMo 12B models to 4B and 8B parameters, respectively, using pruning and distillation [1]. We explore two distinct pruning strategies: (1) depth pruning and (2) joint hidden/attention/MLP (width) pruning, and evaluate the results on common benchmarks from the LM Evaluation Harness [2]. The models are then aligned with NeMo Aligner and tested in instruct-tuned versions. This approach produces a compelling 4B model from Llama 3.1 8B and a state-of-the-art Mistral-NeMo-Minitron-8B (MN-Minitron-8B for brevity) model from Mistral NeMo 12B. We found that with no access to the original data, it is beneficial to slightly fine-tune teacher models on the distillation dataset. We open-source our base model weights on Hugging Face with a permissive license.
Models on Hugging Face: Mistral-NeMo-Minitron-8B-Base | Llama-3.1-Minitron-4B-Width-Base | Llama-3.1-Minitron-4B-Depth-Base
Benchmarks(shots)
Gemma2
Minitron
Llama-3.1-Minitron
Gemma
Mistral
Llama 3.1
MN-Minitron
Mistral NeMo
2B* 4B 4B-Depth 4B-Width 7B 7B 8B 8B 12B-Base 12B-FT
Total Params 2.6B 4.2B 4.5B 4.5B 8.5B 7.3B 8B 8.4B 12.2B 12.2B
Non-Emb. Params 2B 2.6B 3.7B 3.7B 7.7B 7B 7B 7.3B 10.9B 10.9B
Training Tokens 2T 94B 94B 94B 6T 8T 15T 380B - +0.1T
Winogrande(5) 70.9 74.0 72.1 73.5 78 78.5 77.3 80.4 82.2 82.7
Arc_challenge(25) 55.4 50.9 52.6 55.6 61 60.3 57.9 64.4 65.1 62.3
MMLU(5) 51.3 58.6 58.7 60.5 64 64.1 65.3 69.5 69.0 70.1
Hellaswag(10) 73.0 75.0 73.2 76.1 82 83.2 81.8 83.0 85.2 85.3
GSM8k(5) 23.9 24.1 16.8 41.2 50 37.0 48.6 58.5 56.4 55.7
Truthfulqa(0) - 42.9 38.2 42.9 45 42.6 45.0 47.6 49.8 48.3
XLSum en(20%) (3) - 29.5 27.2 28.7 17 4.8 30.0 32.0 33.4 31.9
MBPP(0) 29.0 28.2 30.7 32.4 39 38.8 42.3 43.8 42.6 47.9
HumanEval(n=20)(0) 20.1 23.3 - - 32.0 28.7 24.8 36.2 23.8 23.8
Table 1: Accuracy numbers for our MN-Minitron-8B and Llama-3.1-Minitron-4B models. We compare our models to similarly-sized SoTA open models on a variety of common language modeling benchmarks. All evaluations are conducted by us, except entries marked with * (taken from corresponding papers).
Benchmarks
Gemma
Phi-2
Gemma2 Qwen2 Minitron
Llama-3.1-Minitron
2B 2.7B 2B 1.5B 4B 4B-Depth 4B-Width
Total Params 2.5B 2.7B 2.6B 1.5B 4.2B 4.5B 4.5B
Non-Emb. Params 2B 2.5B 2B 1.3B 2.6B 3.5B 3.7B
Tokens 3T 1.4T 2T 7T 94B 94B 94B
IFEval 40.5 44.0 64.5 39.8 44.8 42.6 52.4
MT-Bench 5.2 4.3 7.7 5.2 5.6 5.6 6.3
ChatRAG* 33.3 37.6 37.5 32.8 41.1 40.1 44.0
BFCL 47.0 23.1 35.6 32.8 64.2 66.8 64.9
Table 2: Accuracy numbers for the aligned Llama-3.1-Minitron models. We compare our models to similarly-sized SoTA open aligned models on a variety of benchmarks. All evaluations are conducted by us. * Denotes results obtained on a representative subset of the benchmark. Best in bold, second underlined. The alignment of MN-Minitron-8B is underway and will be posted once ready.
1 Introduction
LLM providers often train an entire family of models from scratch, each with a different size (number of parameters, e.g. Llama 3.1 8B, 70B, 405B); this is done to aid users targeting different deployment scales, sizes and compute budgets. However, training multiple multi-billion parameter models from scratch is extremely time-, data- and resource-intensive.
Recent work [1] has demonstrated the effectiveness of combining weight pruning with knowledge distillation to significantly reduce the cost of training LLM model families. Here, only the biggest model in the family is trained from scratch; other models are obtained by successively pruning the bigger model(s) and then performing knowledge distillation [3] to recover the accuracy of pruned models.
In this report, we successfully apply the Minitron compression strategy [1] to two state-of-the-art models: Llama 3.1 8B [4] and Mistral NeMo 12B [5], compressing them down to 4B and 8B parameters, respectively. Figure 1 provides a high-level overview of our approach.
While following the original paper [1], we make a key modification: due to lack of access to the original training data, we fine-tune the teacher model on our own dataset before pruning and distillation. We refer to this step as teacher correction. Figure 4 shows that omitting teacher correction causes a data distribution mismatch, negatively impacting distillation.
[…]
Figure 1: High-level overview of our proposed pruning and distillation approach. The total number of tokens used for each step is indicated in parentheses.
2 Methodology
A high-level overview of our approach is illustrated in Figure 1. Here, the teacher model is first lightly finetuned on the target dataset to be used for distillation - we refer to this step as teacher correction. Next, pruning is applied to compress the model, following which distillation is used to recover any lost model accuracy. We refer the reader to the Minitron paper [1] for the full description of the pruning and distillation method.
2.1 Pruning
Weight pruning is a powerful and well-known technique for reducing model size. In this report, we focus on structured pruning, where blocks (or channels) of nonzero elements are removed at once from model weights; examples of structured pruning techniques include neuron, attention head, convolutional filter, and depth pruning [1]. In case of LLMs, as shown in Figure 2, we start the pruning process by first computing the importance of each layer, neuron, head, and embedding dimension. We then sort these importance scores to compute a corresponding importance ranking.
Figure 2: Pruning and distillation process outlined in the original paper [1]. We follow the same approach in this work.
Importance Estimation: We use a purely activation-based importance estimation strategy that simultaneously computes sensitivity information for all the axes we consider (depth, neuron, head, and embedding channel) using a small calibration dataset and only forward propagation passes. We consider depth pruning as a special case and do not combine it with compressing other dimensions.
We compute the importance of each head, neuron and embedding channel by examining the activations produced by the multi-head attention (MHA), multi-layer perceptron (MLP) and LayerNorm layers, respectively. We use a small calibration dataset (1024 samples) for this purpose.
For depth pruning, we consider three distinct metrics for evaluating layer importance: (1) LM validation loss, (2) Block Importance (BI) [6] and (3) accuracy on the downstream task. For loss-based ranking, we simply remove a single or a block of contiguous layers and compute its effect on LM loss; this serves as the “importance” or sensitivity of the layer. BI uses the cosine distance between the input and output of a layer or a block of layers. We notice that BI and LM loss metrics are highly correlated but do not produce the most accurate pruned model on downstream tasks as shown in Figures 8 and 9. We thus evaluate layer importance using the Winogrande benchmark [7].
Model Trimming: As shown in Figure 2, for a given architecture configuration, we first rank the elements of each axis according to the computed importance and perform trimming (reshaping) of the corresponding weight matrices directly. For neuron and head pruning, we trim MLP and MHA layer weights, respectively. In the case of embedding channels, we trim the embedding dimension of the weight matrices in MLP, MHA, and LayerNorm layers. The original approach ([1]) uses Neural Architecture Search (NAS) to find the best architecture; in this work, we skip this step and instead utilize the network architecture-related learnings from the original paper.
2.2 Retraining with Distillation
We use the term retraining to refer to the accuracy recovery process following pruning. In this work, we explore two retraining strategies: (1) conventional training, leveraging ground truth labels, and (2) knowledge distillation using supervision from the unpruned model (teacher). Knowledge Distillation (KD) [3] involves transfer of knowledge from a larger or more complex model called the teacher to a smaller/simpler model called the student. The knowledge transfer is achieved by having the student model mimic the output and/or the intermediate states of the teacher model. In our case, the uncompressed and pruned models correspond to the teacher and student, respectively. For distillation, we follow best practices from our previous work [1] and use forward KL Divergence loss [8] on the teacher and student logits only (following the [3]). This is illustrated in Figure 3.
Figure 3: Overview of Distillation: If the original training data is unavailable, a slight fine-tuning of the teacher model is recommended. Distillation is then performed by minimizing KL divergence on the logits, with the original model as the teacher and the pruned model as the student.
[…]
backgroundcolor=gray!10, roundcorner=10pt
3 Training Details
3.1 Pre-training
Llama 3.1 8B [4] and Mistral NeMo [5] 12B are pretrained on different proprietary datasets, which we do not have access to. According to the Llama 3.1 tech report [4], the 8B model is pretrained on 15T tokens. We start with the corresponding Base models that are openly available online on Hugging Face.
Dataset:
We use the Nemotron-4 curated continued training (CT) dataset [9]  [10] for all our pruning and distillation experiments.
3.2 Pruning
Our simplified pruning recipe is based on the best practices outlined in the Minitron paper [1] and is described in the Methodology section. Specifically, for width pruning, we (1) use l2-norm and mean as the aggregation functions across the batch and sequence dimensions, respectively, and (2) perform single-shot pruning, avoiding iterative approaches. For depth pruning, as described in the Methodology section, we follow the observations from Gromov et al. [11] and drop a continuous subgroup of layers that results in the least accuracy drop on Winogrande [7]. In this work, we skip the lightweight neural architecture search (NAS) phase, and go with a manual architecture configuration for both Llama-3.1-Minitron-4B and MN-Minitron-8B. The architectures we come up with are inspired by the Minitron-4B and Minitron-8B models, and are detailed in Table 3. We now describe the pruning recipes for each of our target compressed models:
LLaMa-3.1-Minitron-4B MN-Minitron
Width Depth 8B
Total params 4.5B 4.5B 8.4B
Non-Emb params 3.7B 3.5B 7.3B
Hidden size 3072 4096 4096
Vocabulary 128256 128256 131072
MLP hidden dim 9216 14336 11520
Depth 32 16 40
Attention groups 8 8 8
Query heads 32 32 32
Head dimension 128 128 128
Table 3: Architecture details of our compressed models.
[…]
backgroundcolor=gray!10, roundcorner=10pt
{mdframed}
Llama-3.1-Minitron-4B-Width:
  • Starting model: Llama 3.1 8B
  • Hidden dimension: 4096 $\rightarrow$ 3072
  • MLP hidden dimension: 14336 $\rightarrow$ 9216
  • Attention heads: unchanged
  • Depth: unchanged
{mdframed}
Llama-3.1-Minitron-4B-Depth:
  • Starting model: Llama 3.1 8B
  • Hidden dimension: unchanged
  • MLP hidden dimension: unchanged
  • Attention heads: unchanged
  • Depth: 32 $\rightarrow$ 16
{mdframed}
MN-Minitron-8B:
  • Starting model: Mistral NeMo 12B
  • Hidden dimension: 5120 $\rightarrow$ 4096
  • MLP hidden dimension: 14336 $\rightarrow$ 11520
  • Attention heads: unchanged
  • Depth: unchanged
3.3 Distillation
Teacher Correction:
[…]
Retraining:
Following the learnings in the Minitron work [1], we opt for logit-only distillation, minimizing the forward KL Divergence [8] loss across the teacher and student probabilities, and ignore the LM cross-entropy loss altogether. Here, the unpruned and pruned models correspond to the teacher and student, respectively. We use the hyperparameters listed in Table 4 during distillation. We use 32 NVIDIA DGX H100 nodes for our training jobs.
Llama-3.1- MN-Minitron
Minitron-4B 8B
Peak learning rate 1e-4 1e-4
Min learning rate 1e-5 4.5e-7
Warm-up steps 40 steps 60 steps
LR decay schedule Cosine Cosine
Global batch size 1152 768
Context length 8192 8192
Total tokens 94B 380B
Table 4: Hyperparameters used during distillation-based retraining.
Figure 4: Training convergence plot for the compressed 8B student model. We compare supervision from the original teacher and the corrected teacher.
3.4 Instruction Tuning
To evaluate the instruction-following capabilities of our distilled models, we perform supervised fine-tuning (SFT) on the Llama-3.1-Minitron 4B models using NeMo-Aligner [12] with the instruction tuning dataset used for Nemotron-4 340B [13]. As shown in Table 2, we evaluate the aligned models for instruction- following and roleplay (IFEval [14] and MT-Bench [15]), RAG QA (ChatRAG-Bench [16]), and function-calling capabilities (BFCL [17]).
4 Analysis
We perform a series of ablation studies to better understand the compression characteristics of these newer models. We report our results in this section.
Width vs Depth Pruning:
Figure 5 shows the training curve of Llama-3.1-Minitron-4B pruned for width vs. depth. We notice that width pruning results in smaller initial loss and consistently outperforms the depth-pruned model, despite both variants having the same number of parameters.
Pruning and Distillation:
Figure  6 demonstrates orthogonal benefits of our proposed approach with pruning and distillation. We compare (1) random weight initialization and distillation, (2) random pruning and distillation, where components are pruned randomly ignoring the importance scores, (3) our proposed pruning with typical cross entropy based LM loss training and (4) our proposed pruning with distillation-based training. We notice that pruning results in a significantly better starting point compared to random initialization, and also that distillation-based training outperforms conventional training methods while requiring significantly fewer training tokens (up to $50\times$50 × in our case).
Figure 5: Convergence of width- and depth-pruned Llama 3.1 8B to 4B models. Width pruning consistently outperforms depth pruning for a given parameter budget.
Figure 6: Training convergence plot for Mistral Nemo 12B compressed model. We compare (a) random initialization with distillation, (b) randomly pruned weights with distillation, (c) pruning with standard LM loss, and (d) our pipeline with pruning and distillation.
Teacher Correction:
We compare two approaches for teacher correction: (1) pruning and distilling the corrected teacher, and (2) pruning the original teacher and distilling from a continuously corrected teacher. The results in Figure 7 suggest that teacher correction doesn’t affect the optimality of pruning, and that distillation from a corrected teacher is crucial. Teacher correction can be performed in parallel with distillation to bridge the gap.
Figure 7: Training convergence plot for Mistral Nemo 12B compressed model. We compare (1) pruning and distilling the corrected teacher with (2) pruning the original teacher and distilling from a continuously corrected teacher.
Depth Pruning Metrics:
when examining how LM validation loss increases as contiguous blocks of layers are removed (Figure  8), we observe that the layers at the beginning and end are the most important. Removing non-contiguous layers can result in even better LM validation loss (the dashed line). However, this observation does not necessarily hold when evaluating downstream task performance. Figure  9 shows that dropping 16 layers selected based on per-layer importance ( [6, 18]) yields a random Winogrande accuracy of 0.5, while removing layers 16 to 31 continuously ( [11]) results in an accuracy of 0.595. The gap holds during distillation-based retraining and we opt for the latter approach.
Figure 8: LM loss value on validation set after removing 1, 2, 8 or 16 contiguous layers with Llama 3.1 8B. For example, the purple line at layer no. 16 indicates the LM loss if we dropped the first 16 layers. Layer no. 17 indicates the LM loss if we leave the first layer intact and drop layers 2 to 17. The dashed line corresponds to LM loss value when removing 16 non-contiguous layers least increasing the loss.
Figure 9: Accuracy on the Winogrande task when removing 16 contiguous layers with Llama 3.1 8B. Layer no. 17 indicates the LM loss if we leave the first layer intact and drop layers 2 to 17. The dashed line corresponds to the accuracy when removing 16 non-contiguous layers least increasing the loss.
5 Evaluation
Benchmarks
following Touvron et al. [19], we evaluate our compressed models on a series of downstream tasks, including MMLU [20], HumanEval [21] for Python code generation, several question-answering datasets for common-sense reasoning: Arc-C [22], HellaSwag [23], TruthfulQA [24] and WinoGrande [7] and XL-Sum English [25] for summarization. We report the 5-shot performance on MMLU, 5-shot on Winogrande, 25-shot on ARC-Challenge, 10-shot on HellaSwag, 0-shot on 20% of XL-Sum and average pass@1 scores for HumanEval and MBPP. For pass@1 scores we use a temperature of 0.2 and nucleus sampling [26] with top-p === 0.95. For instruction-tuned models, we use MT-Bench [15], Instruction-Following Eval (IFEval) [14], ChatRAG-Bench [16], and Berkeley Function Calling Leaderboard (BFCL) [17].
5.1 Base Models
[…]
5.2 Instruct Models
The performance of the instruction-tuned Llama-3.1-Minitron 4B variants is shown in Table 2. We compare the Llama-3.1-Minitron 4B variants to other similarly-sized baselines and notice that our models demonstrate strong instruction-following and roleplay capabilities, only lagging behind Gemma2 in IFEval [14] and MT-Bench [15]. On retrieval based question answering (ChatRAG-Bench [16]) and function-calling (BFCL [17]), Minitron models achieve state-of-the-art performance.
5.3 Runtime Performance Analysis
We optimized the Llama 3.1 8B and Llama-3.1-Minitron 4B variants with NVIDIA TensorRT-LLM, an open-source toolkit for optimized LLM inference.
Figure 10: TensorRT-LLM FP8 throughput comparison for the Llama-3.1-Minitron-4B models with the Llama 3.1 8B model w.r.t. increasing input and output sequence lengths.
Figure 10 shows the throughput in requests per second for the various models in FP8 precision obtained on a single H100 80 GB GPU. Different use cases are represented by increasing input sequence length/output sequence length (ISL/OSL) combinations, at a batch size of 32 and 64 for the 8B-12B models and the 4B models respectively. The smaller memory footprint of the 4B model allows for larger batches. We notice that Llama-3.1-Minitron-4B-Depth is the fastest, achieving an average throughput improvement of $2.7\times$2.7 × over Llama 3.1 8B; the Llama-3.1-Minitron-4B-Width variant achieves an average throughput improvement of $1.8\times$1.8 × over Llama 3.1 8B. Compared to BF16, we notice that FP8 delivers a performance boost of $1.4\times$1.4 ×.
6 Insights
In this Section, we summarize some interesting and surprising observations.
6.0.1 General
  1. 1.
    Teacher correction is crucial for distillation to work optimally on a new, unseen dataset. Fine-tuning the teacher with the dataset used for distillation in this manner yields over a 6% reduction in LM validation loss. Teacher correction doesn’t affect the optimality of pruning and can even be performed in parallel with distillation.
  2. 2.
    In line with the Minitron paper’s observations, we require only 380B tokens to achieve state-of-the-art accuracy post pruning with distillation.
  3. 3.
    For width pruning, we achieve stronger accuracy by retaining attention heads and pruning the other dimensions (MLP intermediate dimension, embedding channels).
6.0.2 Mistral NeMo 12B to MN-Minitron-8B:
  1. 1.
    Our compressed model outperforms the teacher on two benchmarks, GSM8k and HumanEval after pruning and distillation: GSM8k increases from 55.7% to 58.5% and HumanEval increases from 23.8% to 36.2%. This improvement is likely influenced by the dataset. However, retraining is performed using the distillation loss alone.
6.0.3 Llama 3.1 8B to Llama-3.1-Minitron 4B:
  1. 1.
    Width pruning delivers better accuracy with MMLU at 60.5%, while depth pruning yields 58.7%, for Llama-3.1 compression.
  2. 2.
    Reasoning ability is impacted further significantly, with GSM8K accuracy at 41.24% for width and 16.8% for depth.
  3. 3.
    Depth pruning boosts throughput, achieving $2.7\times$2.7 × speedup over Llama-3.1 8B, while width pruning provides $1.7\times$1.7 × speedup.
  4. 4.
    For depth pruning, we observe that dropping contiguous layers from the model is more effective than using non-contiguous, importance-based pruning.
7 Acknowledgments
This work would not have been possible without contributions from many people at NVIDIA. To mention a few:
Foundational Model: Sharath Turuvekere Sreenivas, Saurav Muralidharan, Raviraj Joshi, Marcin Chochowski, Pavlo Molchanov, Mostofa Patwary, Daniel Korzekwa, Ashwath Aithal, Mohammad Shoeybi, Bryan Catanzaro and Jan Kautz
Alignment: Ameya Sunil Mahabaleshwarkar, Hayley Ross, Brandon Rowlett, Oluwatobi Olabiyi, Shizhe Diao and Yoshi Suhara
Datasets: Sanjeev Satheesh, Jupinder Parmar, Shengyang Sun, Jiaqi Zeng, Zhilin Wang, Yi Dong, Zihan Liu, Rajarshi Roy, Wei Ping, Makesh Narsimhan Sreedhar and Oleksii Kuchaiev
TensorRT-LLM: Bobby Chen, James Shen and Chenhan Yu
Hugging Face Support: Ao Tang, Yoshi Suhara and Greg Heinrich