### v3

# LLM Pruning and Distillation in Practice: The Minitron Approach

Abstract
Abstract: Structured pruning with knowledge distillation is a potent combination for obtaining small language models (SLMs) with significantly fewer training tokens and compute resources compared to training from scratch. In this work, we investigate how this strategy can be effectively applied in instances where access to the the original pretraining dataset is restricted. We introduce a new teacher correction phase before distillation which lets the teacher model adjust to our specific data distribution using a lightweight fine-tuning phase. We apply this strategy to compress the Mistral NeMo 12B and Llama 3.1 8B models to 8B and 4B parameters, respectively, using pruning and distillation. We explore two distinct pruning strategies: (1) depth pruning and (2) joint hidden/attention/MLP (width) pruning, and evaluate the results on common benchmarks from the LM Evaluation Harness. The models are then aligned with NeMo Aligner and further tested for instruction following, role-play, math, coding and function calling capabilities. This approach produces the state-of-the-art Mistral-NeMo-Minitron-8B (MN-Minitron-8B for brevity) model from Mistral NeMo 12B, and a compelling 4B model from Llama 3.1 8B. We open-source our base model weights on Hugging Face with a permissive license.
Models on Hugging Face: Mistral-NeMo-Minitron-8B-Base | Llama-3.1-Minitron-4B-Width-Base | Llama-3.1-Minitron-4B-Depth-Base
1 Introduction
LLM providers often train an entire family of models from scratch, each with a different size (number of parameters, e.g. Llama 3.1 with 8B, 70B, and 405B parameters [1]); this is done to aid users targeting different deployment scales, sizes and compute budgets. However, training multiple billion-plus parameter models from scratch is extremely time-, data- and resource-intensive. Recent work has demonstrated the effectiveness of combining weight pruning with knowledge distillation to significantly reduce the cost of training LLM model families [2]. Here, only the biggest model in the family is trained from scratch; other models are obtained by successively pruning the bigger model(s) and then performing knowledge distillation [3] to recover the accuracy of pruned models. While highly effective, this line of work assumes access to the original pretraining dataset for the distillation phase. With a growing number of frontier LLMs (including open ones) being trained on private, proprietary datasets [1, 4], this assumption often fails to hold.
Figure 1: High-level overview of our proposed pruning and distillation approach. The total number of tokens used for each step is indicated in parentheses.
Benchmarks (shots)
Gemma2
Minitron
Llama-3.1-Minitron
Gemma
Mistral
Llama 3.1
MN-Minitron
Mistral NeMo
2B* 4B 4B-Depth 4B-Width 7B 7B 8B 8B 12B-Base 12B-FT
Total Params 2.6B 4.2B 4.5B 4.5B 8.5B 7.3B 8B 8.4B 12.2B 12.2B
Non-Emb. Params 2B 2.6B 3.7B 3.7B 7.7B 7B 7B 7.3B 10.9B 10.9B
Training Tokens 2T 94B 94B 94B 6T 8T 15T 380B - +0.1T
Winogrande(5) 70.9 74.0 72.1 73.5 78 78.5 77.3 80.4 82.2 82.7
Arc_challenge(25) 55.4 50.9 52.6 55.6 61 60.3 57.9 64.4 65.1 62.3
MMLU(5) 51.3 58.6 58.7 60.5 64 64.1 65.3 69.5 69.0 70.1
Hellaswag(10) 73.0 75.0 73.2 76.1 82 83.2 81.8 83.0 85.2 85.3
GSM8k(5) 23.9 24.1 16.8 41.2 50 37.0 48.6 58.5 56.4 55.7
Truthfulqa(0) - 42.9 38.2 42.9 45 42.6 45.0 47.6 49.8 48.3
XLSum en(20%) (3) - 29.5 27.2 28.7 17 4.8 30.0 32.0 33.4 31.9
MBPP(0) 29.0 28.2 30.7 32.4 39 38.8 42.3 43.8 42.6 47.9
HumanEval(n=20)(0) 20.1 23.3 - - 32.0 28.7 24.8 36.2 23.8 23.8
Table 1: Accuracy numbers for our MN-Minitron-8B and Llama 3.1-Minitron-4B models. We compare our models to similarly-sized SoTA open models on a variety of common language modeling benchmarks. All evaluations are conducted by us, except entries marked with * (taken from corresponding papers).
Benchmarks (shots)
Phi-2
Gemma2 Qwen2 Minitron
Llama-3.1-Minitron
LLama 3.1
MN-Minitron
2.7B 2B 1.5B 4B 4B-Depth 4B-Width 8B 8B
MT-Bench (GPT4-Turbo) 5.14 7.44 5.49 6.46 6.19 6.88 7.78 7.86
MMLU (5) 56.8 56.9 55.6 59.3 61.21 59.89 69.4* 70.4
GSM8K (0) 19.9 52.2 27.2 65.1 71.11 79.76 83.8 87.1
GPQA (0) 28.8 25.9 28.1 29.5 32.59 30.36 30.4* 31.5
HumanEval (0) 47.6* 45.1 47.0* 39.6 42.7 47.0 72.6 71.3
MBPP (0) 55.0* 50.4 51.9* 57.4 60.3 65.1 72.8* 72.5
IFEval 44.0 64.5 39.8 75.3 66.77 79.54 80.4* 84.4
BFCLv2 (Live) 38.7 40.2 39.9 53.1 55.89 55.0 44.3 67.6
Table 2: Accuracy numbers for instruction tuned models on a variety of benchmarks. All evaluations are conducted by us, except entries marked with * (taken from corresponding papers). Best of each section in bold. For IFEval, we report the average of prompt and instruction across loose and strict evaluations. For BFCLv2, we report live accuracy only.
In this work, we adapt the original Minitron compression recipe [2] along two directions: (1) we introduce a new teacher correction phase for adapting the teacher (unpruned) model to our own data distribution, thus removing any need to access the original pretraining dataset, and (2) we introduce a new and more effective downstream task-based saliency criteria for depth pruning. We successfully apply our updated compression strategy to two state-of-the-art models: Llama 3.1 8B [1] and Mistral NeMo 12B [5], compressing them down to 4B and 8B parameters, respectively. For Llama 3.1 8B, we produce two distinct compressed models in the 4B parameter range: (1) Llama 3.1-Minitron-4B-Width (pruning only the width axes), and (2) Llama 3.1-Minitron-4B-Depth (pruning depth only). Figure 1 provides a high-level overview of our approach.
Tables 1 and 2 provide a summary of our results: our compression strategy yield a state-of-the-art 8B model (MN-Minitron-8B) which outperforms all similarly-sized models across the board on common language modeling benchmarks. Our Llama 3.1-Minitron-4B models (both depth and width-pruned variants) also exhibit strong accuracy compared to the teacher Llama 3.1 8B model and the previous-generation Minitron-4B model [2]; among the two variants, the width-pruned variant achieves better overall accuracy than the depth-pruned one. In terms of runtime inference performance measured using TensorRT-LLM, the Llama 3.1-Minitron-4B models provide an average speedup of 2.7$\times$ and 1.8$\times$ for the depth and width pruned variants, respectively, compared to the original Llama 3.1 8B model.
2 Methodology
A high-level overview of our approach is illustrated in Figure 1. Here, the teacher model undergoes a lightweight adjustment phase on the target dataset to be used for distillation - we refer to this step as teacher correction. Next, pruning is applied to compress the model, following which distillation is used to recover model accuracy.
Figure 2: Overview of distillation: if/when the original training data is unavailable, a lightweight fine-tuning of the original model on the distillation dataset is recommended, to be used as a teacher. Distillation is then performed by minimizing KL divergence on the logits of the teacher and the pruned student model.
2.1 Teacher Correction
Distillation is an effective technique to condense knowledge from a more accurate teacher model to improve a less accurate student model  [3]  [2]. Typically, knowledge is distilled using the same dataset the teacher model was trained on. In cases where access to the original training data is restricted, we notice from our experiments that the teacher model provides sub-optimal guidance if a different dataset is used to distill the knowledge. We hypothesize this is due to the change in distribution of sub-word tokens across the original dataset the teacher model was trained on vs. the dataset being distilled on. To this end, we propose a novel teacher correction phase (illustrated in Figure 2), where we perform a lightweight ($\sim$100B tokens) fine-tuning of the teacher model to adapt to the new distillation dataset. We demonstrate in our experimental evaluation (Figure 5 in particular) that this procedure significantly improves the guidance resulting in a more accurate student model. We also explore correcting the teacher in parallel to distillation, and demonstrate that this performs on par with using guidance from a fully corrected teacher.
Figure 3: Pruning and distillation process outlined in the original paper [2]. We follow the same approach in this work.
2.2 Pruning
Weight pruning is a powerful and well-known technique for reducing model size. In this work, we focus on structured pruning, where blocks (or channels) of nonzero elements are removed at once from model weights; examples of structured pruning techniques include neuron, attention head, convolutional filter, and depth pruning [6, 7, 8, 9]. We follow the pruning recipe outlined in Minitron [2]: as shown in Figure 3, we start the pruning process by first computing the importance of each layer, neuron, head, and embedding dimension. We then sort these importance scores to compute a corresponding importance ranking.
Importance Estimation We use a purely activation-based importance estimation strategy that simultaneously computes sensitivity information for all the axes we consider (depth, neuron, head, and embedding channel) using a small calibration dataset and only forward propagation passes. We consider depth pruning as a special case and do not combine it with compressing other dimensions. We compute the importance of each head, neuron and embedding channel by examining the activations produced by the multi-head attention (MHA), multi-layer perceptron (MLP) and LayerNorm layers, respectively. We use a small calibration dataset (1024 samples) drawn randomly from the full dataset for this purpose.
Layer Importance For depth pruning, we consider two distinct metrics for evaluating layer importance: (1) LM validation loss/PPL, and (2) accuracy on the downstream task. We do not consider the Block Importance (BI) metric [8] as it was recently shown to under-perform the validation loss/PPL metric [2]. For ranking, we simply remove a single or a block of contiguous layers and compute its effect on each metric; this serves as the “importance” or sensitivity of the layer/layer block. Based on our empirical analysis (see Figures 9 and 9), we use the Winogrande metric [10] to prune sets of contiguous layers. This pruning strategy evolved from two important observations: (1) LM validation loss/PPL-based layer importance fails to produce the most accurate pruned model(s) on downstream tasks, and (2) dropping contiguous layers is better than individual, as also observed in Gromov et al. [11].
Model Trimming Following Minitron [2], for a given architecture configuration, we first rank the elements of each axis according to the computed importance and perform trimming of the corresponding weight matrices directly. For neuron and head pruning, we trim MLP and MHA layer weights, respectively. In the case of embedding channels, we trim the embedding dimension of the weight matrices in MLP, MHA, and LayerNorm layers. The original approach uses Neural Architecture Search (NAS) to find the best architecture; in this work, we skip this step and instead utilize the network architecture-related learnings from the original paper.
2.3 Retraining with Distillation
We use the term retraining to refer to the accuracy recovery process post pruning. In this work, we explore two retraining strategies: (1) conventional training, leveraging ground truth labels, and (2) knowledge distillation using supervision from the unpruned model (teacher). Knowledge Distillation (KD) [3] involves transfer of knowledge from a larger or more complex model called the teacher to a smaller/simpler model called the student. The knowledge transfer is achieved by having the student model mimic the output and/or the intermediate states of the teacher model. In our case, the uncompressed and pruned models correspond to the teacher and student, respectively. Following the best practices outlined in the Minitron work [2], we use forward KL Divergence loss [12] on the teacher and student logits only; this is illustrated in Figure 2.
3 Training Details
3.1 Pre-training
Llama 3.1 8B [1] and Mistral NeMo 12B [5] are pretrained on different proprietary datasets, which we do not have access to. According to the Llama 3.1 tech report [1], the 8B model is pretrained on 15T tokens. We start with the corresponding Base models that are openly available on Hugging Face.
Dataset
We use the Nemotron-4 curated continued training (CT) dataset [13]  [14] for all our pruning and distillation experiments.
3.2 Teacher Correction
Using the original Mistral NeMo 12B or Llama 3.1 8B models directly as a teacher performs sub-optimally on our dataset. To counter this, we apply teacher correction, as described in the previous section, to both models with $\sim 100B$ tokens. Since the goal is to adapt the teacher model to the distillation dataset, we use 120 steps of warm-up and low learning rates: one-fifth the peak learning rate, identical batch size, minimum learning rate and decay schedule the original model was trained on. We notice that the correction process has a minor effect on the teacher model’s accuracy on downstream tasks, with some tasks improving and some degrading as shown in Table 1. We hypothesize this to be an artifact of the dataset used for fine-tuning. Optimizing this process further by using fewer than $\sim$100B tokens, lighter fine-tuning such as LoRA [15] or tuning layer normalization [16] parameters alone would be an interesting topic for future work.
3.3 Pruning
Our pruning recipe is based on the best practices outlined in the Minitron paper [2], as described in the previous section. Specifically, for width pruning, we (1) use l2-norm and mean as the aggregation functions across the batch and sequence dimensions, respectively, and (2) perform single-shot pruning, avoiding iterative approaches. For depth pruning, we follow the observations from Gromov et al. [11] and drop a continuous subgroup of layers that results in the least accuracy drop on Winogrande [10]. In this work, we skip the lightweight neural architecture search (NAS) phase, and go with a manual architecture configuration for both Llama 3.1-Minitron-4B and MN-Minitron-8B. The architectures we come up with are inspired by the Minitron-4B and Minitron-8B models [2], and are detailed in Table 3. We provide the pruning recipes for each of our target compressed models below:
LLaMa-3.1-Minitron MN-Minitron
4B-Width 4B-Depth 8B
Total params 4.5B 4.5B 8.4B
Non-Emb params 3.7B 3.5B 7.3B
Hidden size 3072 4096 4096
Vocabulary 128256 128256 131072
MLP hidden dim 9216 14336 11520
Depth 32 16 40
Attention groups 8 8 8
Query heads 32 32 32
Head dimension 128 128 128
Table 3: Architecture details of our compressed models.
Llama-3.1-Minitron MN-Minitron
Peak learning rate 1e-4 1e-4
Min learning rate 1e-5 4.5e-7
Warm-up steps 40 steps 60 steps
LR decay schedule Cosine Cosine
Global batch size 1152 768
Context length 8192 8192
Total tokens 94B 380B
Table 4: Hyperparameters used during distillation-based retraining.
[…]
backgroundcolor=gray!10, roundcorner=10pt {mdframed} Llama-3.1-Minitron-4B-Width:
  • Starting model: Llama 3.1 8B
  • Hidden dimension: 4096 $\rightarrow$ 3072
  • MLP hidden dimension: 14336 $\rightarrow$ 9216
  • Attention heads: unchanged
  • Depth: unchanged
{mdframed}
Llama-3.1-Minitron-4B-Depth:
  • Starting model: Llama 3.1 8B
  • Hidden dimension: unchanged
  • MLP hidden dimension: unchanged
  • Attention heads: unchanged
  • Depth: 32 $\rightarrow$ 16
{mdframed}
MN-Minitron-8B:
  • Starting model: Mistral NeMo 12B
  • Hidden dimension: 5120 $\rightarrow$ 4096
  • MLP hidden dimension: 14336 $\rightarrow$ 11520
  • Attention heads: unchanged
  • Depth: unchanged
3.4 Distillation
We opt for logit-only distillation, minimizing the forward KL Divergence [12] loss across the teacher and student probabilities, and ignore the LM cross-entropy loss altogether. Here, the unpruned and pruned models correspond to the teacher and student, respectively. We use the hyperparameters listed in Table 4 during distillation. We use 32 NVIDIA DGX H100 nodes for our training jobs.
Figure 4: Training convergence plot for the MN-Minitron-8B student model. We compare supervision from the original teacher and the corrected teacher.
Figure 5: Training convergence plot for the MN-Minitron-8B student model. We compare (1) pruning and distilling the corrected teacher with (2) pruning the original (uncorrected) teacher and distilling from a continuously corrected teacher. We notice that teacher correction can be performed in parallel with distillation.
3.5 Instruction Tuning
To evaluate the instruction-following capabilities of our distilled models, we perform alignment using NeMo-Aligner [17]. We follow the same recipe for all our models by first applying math and code supervised fine-tuning (SFT) followed by instruction SFT and then two rounds of Reward-aware Preference Optimization (RPO) [18].
4 Analysis
We perform a series of ablation studies to better understand the effects of distillation, teacher correction, and our new depth-pruning saliency metric. We report our findings in this section.
Teacher Correction
We first compare the effects of teacher correction on the MN-Minitron-8B model in Figure 5; here, we notice the clear benefits of performing teacher correction w.r.t. distilling directly from an uncorrected teacher. Next, we compare two approaches for teacher correction: (1) pruning and distilling the corrected teacher, and (2) pruning the original (uncorrected) teacher and distilling from a continuously corrected teacher. The results in Figure 5 suggest that teacher correction can be performed in parallel with distillation to recover accuracy of the pruned student model.
Pruning and Distillation
Figure 7 demonstrates the orthogonal benefits of pruning and distillation over random initialization and conventional fine-tuning, respectively. We compare (1) random weight initialization and distillation, (2) random pruning and distillation, where weights are pruned randomly ignoring the importance scores, (3) our proposed pruning with typical cross entropy based LM loss training and (4) our proposed pruning with distillation-based retraining. We notice that pruning results in a significantly better starting point compared to random initialization, and distillation-based training outperforms conventional training methods. Overall, our approach requires significantly fewer training tokens (up to 40×40$\times$40 ×; 380B instead of 15T tokens) to produce the state-of-the-art MN-Minitron-8B model.
Width vs. Depth Pruning
Figure 7 shows the training curve of Llama 3.1-Minitron-4B pruned for width vs. depth. We notice that width pruning results in a lower initial loss and consistently outperforms the depth-pruned model, despite both variants having the same number of parameters.
Figure 6: Training convergence plot for the MN-Minitron-8B model. We compare (a) random initialization with distillation, (b) randomly pruned weights with distillation, (c) pruning with standard LM loss, and (d) our pipeline with pruning and distillation. This plot shows the benefits of pruning and distillation over random initialization and conventional finetuning, respectively.
Figure 7: Convergence plots for the width-pruned and depth-pruned versions of Llama 3.1 8B to 4B compressed models. Width pruning consistently outperforms depth pruning for a given parameter budget.
Depth Pruning Metrics
By examining how LM validation loss increases as contiguous blocks of layers are removed (Figure  9), we observe that the layers at the beginning and end are the most important. The figure indicates that removing non-contiguous layers can result in even better LM validation loss (the dashed line). However, we notice this observation does not necessarily hold when evaluating downstream task performance: specifically, Figure 9 shows that dropping 16 layers selected based on per-layer importance [8, 19] yields a random Winogrande accuracy of 0.5, while removing layers 16 to 31 continuously [11] results in an accuracy of 0.595. The gap holds during distillation-based retraining and we opt for the latter approach in this work.
Figure 8: LM loss value on validation set after removing 1, 2, 8 or 16 contiguous layers from Llama 3.1 8B. The purple line at layer no. 16 indicates the LM loss if we dropped the first 16 layers. Layer no. 17 indicates the LM loss if we leave the first layer intact and drop layers 2 to 17. The dashed line corresponds to LM loss value when removing 16 non-contiguous layers least increasing the loss.
Figure 9: Accuracy on the Winogrande task when removing 16 contiguous layers from Llama 3.1 8B. Layer no. 17 indicates the accuracy if we leave the first layer intact and drop layers 2 to 17. The dashed line corresponds to the accuracy when removing 16 non-contiguous layers that increasing the loss by the least amount.
5 Evaluation
Benchmarks
following Llama [20], we evaluate our compressed base and aligned models on a series of downstream tasks, namely MMLU [21], HumanEval [22] for Python code generation, MBPP [23] and GSM8K [24]. We also evaluate the base models on several question-answering datasets for common-sense reasoning: Arc-C [25], HellaSwag [26], TruthfulQA [27], WinoGrande [10], and XL-Sum English [28] for summarization. The instruction tuned models are further evaluated for question-answering, function calling, instruction following and multiturn conversations on GPQA [29], BFCL [30], IFEval [31] and MT-Bench (GPT4-Turbo) [32], respectively. Note that this MT-Bench is a corrected version of the original MT-Bench [33].
For base models, accuracy is reported with the following evaluations settings: 5-shot on MMLU, 5-shot on Winogrande, 25-shot on ARC-Challenge, 10-shot on HellaSwag, 0-shot on 20% of XL-Sum and average pass@1 scores for HumanEval and MBPP. For pass@1 scores we use a temperature of 0.2 and nucleus sampling [34] with top-p $=$ 0.95. For aligned models we use 0 shot and greedy sampling if applicable.
5.1 Base Models
Base model evaluation results are shown in Table 1. Compared to similarly-sized models, MN-Minitron-8B demonstrates superior accuracy across the board, outperforming the recent Llama 3.1 8B model using 40$\times$ fewer training tokens (380B vs. 15T). Similarly, the Llama 3.1-Minitron-4B models perform favorably compared to the teacher Llama 3.1 8B model using 150×150$\times$150 × fewer training tokens (94B vs. 15T); our pruned Llama models also outperform the original Minitron 4B model [2]. We note from Table 1 that the width-pruned Llama variant outperforms the depth-pruned one. These results clearly demonstrate the advantages of our methodology: state-of-the-art accuracy coupled with an order of magnitude improvement in training efficiency.
5.2 Instruct Models
The accuracy of the instruction-tuned model variants are shown in Table 2. Our aligned models outperform similarly sized variants on most evaluated benchmarks with the exception of HumanEval [35] and MBPP [23]. Additionally, Llama 3.1-Minitron-4B lags behind Gemma2 on MT-Bench [33]. Nevertheless, our aligned models are consistently better on MMLU [21], GSM8K [24], GPQA [29], IFEval [31] and BFCLv2 [30]. This demonstrates the strong capabilities of our model.
5.3 Runtime Performance Analysis
To evaluate runtime performance, we optimize the Llama 3.1 8B and Llama 3.1-Minitron-4B variants with NVIDIA TensorRT-LLM, an open-source toolkit for optimized LLM inference.
Figure 10: TensorRT-LLM FP8 throughput comparison for the Llama 3.1-Minitron-4B models with the Llama 3.1 8B model w.r.t. increasing input and output sequence lengths.
Figure 10 shows the throughput in requests per second for the various models in FP8 precision obtained on a single H100 80 GB GPU. Different use cases are represented by increasing input sequence length/output sequence length (ISL/OSL) combinations, at a batch size of 32 and 64 for the 8B-12B models and the 4B models respectively. The smaller memory footprint of the 4B model allows for larger batches. We notice that Llama 3.1-Minitron-4B (Depth) is fastest, achieving an average throughput improvement of 2.7×2.7$\times$2.7 × over Llama 3.1 8B; the width-pruned variant achieves an average throughput improvement of 1.8×1.8$\times$1.8 × over Llama 3.1 8B. Compared to BF16, we notice that FP8 delivers a performance boost of 1.4×1.4$\times$1.4 ×.
6 Insights
In this section, we summarize some interesting and surprising observations based on our evaluation.
General
  1. 1.
    Teacher correction is crucial for distillation to work optimally on a new, unseen dataset. Fine-tuning the teacher with the dataset used for distillation in this manner yields over a 6% reduction in LM validation loss. Teacher correction doesn’t affect the optimality of pruning and can even be performed in parallel with distillation.
  2. 2.
    In line with the Minitron paper’s observations, we require a order of magnitude fewer tokens (380B vs 15T) to achieve state-of-the-art accuracy post pruning with distillation.
  3. 3.
    For width pruning, we achieve stronger accuracy by retaining attention heads and pruning the other dimensions (MLP intermediate dimension, embedding channels).
Mistral NeMo 12B to MN-Minitron-8B
  1. 1.
    Our compressed model outperforms the teacher on two benchmarks, GSM8k and HumanEval after pruning and distillation: GSM8k increases from 55.7% to 58.5% and HumanEval increases from 23.8% to 36.2%. This improvement is likely influenced by the dataset. However, retraining is performed using the distillation loss alone.
Llama 3.1 8B to Llama 3.1-Minitron-4B
  1. 1.
    Width pruning delivers better accuracy with MMLU at 60.5%, while depth pruning yields 58.7%, for Llama 3.1 compression.
  2. 2.
    Reasoning ability for base variants appears to be impacted significantly for the depth pruned version, with GSM8K accuracy at 16.8% compared to 41.24% for the width pruned version. However, the gap reduces with instruct tuning.
  3. 3.
    Depth pruning boosts throughput, achieving 2.7×2.7$\times$2.7 × speedup over Llama-3.1 8B, while width pruning provides 1.7×1.7$\times$1.7 × speedup.
  4. 4.
    For depth pruning, we observe that dropping contiguous layers from the model is more effective than using non-contiguous, importance-based pruning.
7 Acknowledgments
This work would not have been possible without contributions from many people at NVIDIA. To mention a few:
Foundational Model: Sharath Turuvekere Sreenivas, Saurav Muralidharan, Raviraj Joshi, Marcin Chochowski, Pavlo Molchanov, Mostofa Patwary, Daniel Korzekwa, Ashwath Aithal, Mohammad Shoeybi, Bryan Catanzaro and Jan Kautz
Alignment: Ameya Sunil Mahabaleshwarkar, Hayley Ross, Brandon Rowlett, Oluwatobi Olabiyi, Shizhe Diao and Yoshi Suhara
Datasets: Sanjeev Satheesh, Jupinder Parmar, Shengyang Sun, Jiaqi Zeng, Zhilin Wang, Yi Dong, Zihan Liu, Rajarshi Roy, Wei Ping, Makesh Narsimhan Sreedhar and Oleksii Kuchaiev
TensorRT-LLM: Bobby Chen, James Shen and Chenhan Yu
Hugging Face Support: Ao Tang, Yoshi Suhara and Greg Heinrich